# OVSeg3R: Learn Open-vocabulary Instance Segmentation from 2D via 3D Reconstruction

### v1

@@ 1 Introduction
Recent advances in 3D reconstruction (Murai et al., 2025; Wang et al., 2024; 2025) have made scene geometry capturing accessible. Yet downstream tasks such as manipulation (Liu et al., 2024c; Black et al., 2024), navigation (Song et al., 2025), and augmented reality (AR) require recognizing objects with instance-level identities and locations. Such demand has driven a growing interest in 3D instance segmentation and its open-vocabulary generalization.
Despite major progress in 2D segmentation (Li et al., 2023), where open-vocabulary capabilities already meet most downstream demands (Ren et al., 2024), 3D instance segmentation capabilities remain limited. This limitation persists even though the cost of acquiring 3D scenes has been greatly reduced by the remarkable progress of 3D reconstruction techniques, as the cost of acquiring 3D annotations is still expensive
Therefore, how to leverage diverse 3D scenes provided by 3D reconstruction models and robust 2D masks provided by well-studied 2D perception models to enhance 3D instance segmentation has become an important research topic. Some approaches (Takmaz et al., 2023) train 3D segmentation models solely to produce class-agnostic 3D masks, which are then projected onto 2D to retrieve category information from 2D foundation models (Radford et al., 2021). While this strategy can effectively exploit the strong classification ability of 2D models, it remains limited by the scarcity of 3D annotations, making it difficult to generate reliable 3D masks for unseen objects and often resulting in missed detections. Other methods (Yang et al., 2023) project each view’s 2D segmentation results (Kirillov et al., 2023) into 3D space using the 2D pixel to 3D point correspondences provided by a 3D reconstruction model, and merge the projected masks that belong to the same instance through heuristic strategies. Although this line of work benefits from the classification and segmentation strengths of 2D models, the hand-crafted merging process is error-prone, rendering these methods fragile and performance-constrained. While these methods demonstrate open-vocabulary potential, they over-rely on 2D outputs, leaving native 3D perception ability underdeveloped, which is crucial for advancing 3D understanding (Peng et al., 2023; Ding et al., 2023). Although some works try to distill from 2D models, they need to train 3D Gaussian (Kerbl et al., 2023) to create 2D and 3D correspondences, which is redundant since reconstruction inherently provides this. Moreover, these methods either require per-scene optimization (Lyu et al., 2024; Ye et al., 2024) or high-quality point cloud for Gaussian initialization (Cao et al., 2025), limiting their practicality.
@@
To address the 3D annotation challenges, in this work, we propose a novel training scheme, called OVSeg3R. As shown in Fig 1, instead of relying on manually adjusted 3D scenes from non-routine sensors, OVSeg3R directly leverages a modern 3D reconstruction model such as (Murai et al., 2025) or (Wang et al., 2025) to provide point cloud inputs. This not only substantially reduces the cost of acquiring 3D scenes, but also naturally introduces noise to the inputs, aligning with application scenarios, where user-provided inputs are typically videos or low-quality reconstructions captured by handheld sensors. For annotations, instead of manually annotating 3D masks for the reconstructed scene, the inherent 2D pixel-to-3D point correspondence provided by 3D reconstruction enables us to lift the 2D masks generated by an open-vocabulary 2D segmention model (Ren et al., 2024) from each view into 3D space to obtain view-level annotations. However, although 2D masks from different views may correspond to the same object, they are estimated independently and lack cross-view associations (Yang et al., 2023). Directly concatenating them to a scene-level annotation would introduce many duplicate annotations. Conversely, since each view covers only part of the reconstructed point cloud, supervising scene-level predictions with each view’s partial annotations alone would incorrectly penalize predictions outside the view. To mitigate this issue, we propose a View-wise Instance Partition (VIP) algorithm. According to the visibility of each scene-level mask prediction’s corresponding object query in different views, VIP assigns the mask predictions to their corresponding views. Then, for each mask prediction, VIP further truncates it to retain only the region visible within its belonging view. The resulting view-level predictions enable supervision with view-specific annotations, which eliminates annotation duplication and avoids incorrect penalization, thus improving training stability.
@@
Moreover, to improve efficiency, mainstream 3D instance segmentation models (Kolodiazhnyi et al., 2024b; Qu et al., 2025) normally leverage superpoints (Landrieu & Simonovsky, 2018), first over-segmenting the input point cloud into superpoints and then performing instance segmentation at the superpoint level. However, in OVSeg3R, 3D reconstruction results are often over-smoothed (Yang et al., 2024), leading to the loss of geometric details. As a result, constructing superpoints purely based on geometric continuity, as in the previous method, may cause objects that are not geometrically salient, such as paintings, being included in a superpoint covering large planar regions, such as walls, as shown in Fig 3 (b). This will inevitably lead to inaccurate segmentation. To mitigate this issue, we propose the 2D Instance Boundary-aware Superpoint (IBSp), IBSp incorporates 2D instance masks into the construction of superpoints, avoids the erroneous clustering of points from different instances into the same superpoint, further stabilizes the training process.
@@ 2 Related Works
Closed-vocabulary 3D Instance Segmentation. Early 3D instance segmentation methods fall into two lines: proposal-based methods (Yang et al., 2019; Hou et al., 2019; Yi et al., 2019; Engelmann et al., 2020; Kolodiazhnyi et al., 2024a), which first detect objects and then refine 3D masks within the predicted bounding boxes, and grouping-based methods (Liang et al., 2021; Chen et al., 2021; Vu et al., 2022; Jiang et al., 2020b; Wang et al., 2019; Jiang et al., 2020a; Zhang & Wonka, 2021), which aggregate points via voting in feature or geometric space. Following the success of Detection Transformers (DETR) (Carion et al., 2020; Liu et al., 2021; Li et al., 2022; Zhang et al., 2022) in 2D, recent works (Sun et al., 2023; Schult et al., 2023; Lai et al., 2023; Kolodiazhnyi et al., 2024b; Jain et al., 2024) adopt DETR-like architectures for 3D instance detection and segmentation. Notably, SegDINO3D (Qu et al., 2025) proposes to leverage high-quality image- and object-level features from well-studied 2D models (Ren et al., 2024; Liu et al., 2024b) to support data-hungry 3D models, achieving substantial performance gains.
Open-vocabulary 3D Instance Segmentation. Motivated by the rapid progress in 2D open-vocabulary perception, most methods attempt to obtain open-vocabulary 3D perception outputs by relying on the outputs of 2D models. Generally, the mainstream methods can be categorized into two types. OpenMASK3D (Takmaz et al., 2023) first proposes to generate class-agnostic 3D segmentation results using 3D models (Schult et al., 2023). Then project each 3D result to 2D to obtain the corresponding object’s category using 2D foundation models (Radford et al., 2021), thereby constructing an open-vocabulary 3D segmentor. While this approach achieves remarkable results and has inspired a series of subsequent works (Nguyen et al., 2024; Boudjoghra et al., 2025; Nguyen et al., 2025), it only provides classification capability for novel categories. Limited by the scarcity of 3D segmentation data, such methods often fail to segment objects that are unseen during training. SAM3D (Yang et al., 2023) is the first to lift 2D segmentation results from SAM for each frame to 3D and then merge those belonging to the same instance in 3D using a proposed heuristic algorithm. Motivated by the promising performance, many subsequent works (Yin et al., 2024; Lu et al., 2023; Zhao et al., 2025; Xu et al., 2025) have focused on improving this merging process. While these algorithms can reuse both the detection and classification capabilities of well-studied 2D models, the heuristics are fundamentally rule-based, lack generalization, and fail to handle numerous corner cases encountered in practical scenarios.
@@ 3 Method
To validate the effectiveness of the proposed training scheme, we adopt SegDINO3D (Qu et al., 2025), a recent state-of-the-art method, as our baseline. To further satisfy the requirement of open-vocabulary, we extend the classification part of SegDINO3D to the similarity calculation between object features and the text features, yielding SegDINO3D-VL. In this section, we will first describe the training of SegDINO3D-VL with OVSeg3R, including the preparation of data, construction of view-wise annotation, and obtaining predictions from SegDINO3D-VL. After that, we will describe in detail our designs for stable training.
@@ 3.1.1 Foundation Preparation
Given a video with $V$ views $\mathbf{I}\in\mathbb{R}^{V\times H\times W\times 3}$, where $H$ and $W$ are height and width, OVSeg3R feeds it to a 3D reconstructor and a 2D instance segmentator, instantiated as MASt3R-SLAM (Murai et al., 2025) and DINO-X (Ren et al., 2024) respectively by default, to prepare the foundation data.
@@ 3.1.2 Input and View-wise Annotation Preparation
2D Feature for Each 3D Point. Following SegDINO3D (Qu et al., 2025), we need to sample 2D image-level feature $\mathbf{F^{2D}}\in\mathbb{R}^{N\times C}$ for each 3D point. However, since the 2D-3D correspondences $\mathcal{C}$ are already available, re-projecting every 3D point to all views to identify a representative view, as required in SegDINO3D, becomes unnecessary. As described in ① of Fig 2, based on each 3D point’s corresponding view index and the sampling location provided by $\mathcal{C}$, we sample 2D features from 2D image feature maps $\mathbf{X}$ for the point through bilinear interpolation. For the sampling of $\mathbf{F}^{\mathbf{2D}}_{i}\in\mathbb{R}^{C}$ for the $i$-th 3D point, the process can be formulated as
@@
Text Prompt Feature. Following previous methods (Liu et al., 2024b), we need to prepare the text features for each class name in $s$ for instance classification (③ in Fig 2). To enable contrastive learning and batch-friendly training, we randomly sample T|s|T-|s| additional class names not present in $s$ as negative classes, padding $s$ to a fixed size $T$. The padded $s$ is then concatenated into a string (e.g., ‘book . sofa .’) and encoded with a text encoder, to produce text features $\mathbf{T}\in\mathbb{R}^{T\times C}$.
@@ 3.2 View-wise Instance Partition
Compatibility with Scene-level 3D Instance Segmentation Supervision. It is worth noting that VIP is fully compatible with standard 3D instance segmentation supervision. For datasets that already provide scene-level annotations, there is no need to partition predictions to individual views for supervision; we can directly supervise the scene-level predictions using the scene-level annotations. This compatibility allows us to mix traditional costly annotated datasets, which provide manually adjusted reconstructed scenes and corresponding human annotations, with datasets that only have the raw videos, thereby enhancing the flexibility of the OVSeg3R training scheme. Moreover, for scene-level supervision, instead of performing a global matching between predictions and annotations via Hungarian matching (Carion et al., 2020), we follow the previous method (Kolodiazhnyi et al., 2024b; Qu et al., 2025) to sparsify the matching based on the relationship between the superpoints used to initialize object queries and the ground-truth masks. Specifically, a prediction can be matched to a ground-truth annotation only if the superpoint used to initialize its object query lies within that annotation’s mask. This design strengthens the connection between the superpoint used for object query initialization and the final mask prediction, thereby consolidating the foundation of VIP from the perspective of supervision and improving its reliability.
@@
Analysis. Current methods typically employ the Felzenszwalb (Felzenszwalb & Huttenlocher, 2004), a graph-based segmentation algorithm, to over-segment large point sets into compact superpoints, reducing the following computational overhead. The algorithm constructs a graph by connecting each 3D point to its K-Nearest Neighbors (KNN). Subsequently, adjacent points in the graph that satisfy geometric continuity constraints will be clustered into the same superpoint.
@@ 4.1 Datasets
We validate the effectiveness of OVSeg3R based on ScanNetv2 (Dai et al., 2017) and ScanNet200 (Rozenberszki et al., 2022). They share the same 1,513 scenes, with 1,201 used for training and the remaining 312 for evaluation. The difference is that ScanNetv2 provides human-annotated instance masks for only 20 classes, whereas ScanNet200 extends the annotations to 200 classes. For each scene, both manually refined high-quality point clouds and the corresponding raw RGB videos are available. We reconstruct the raw videos using MASt3R-SLAM (Murai et al., 2025) and VGGT (Wang et al., 2025), resulting in reconstructed versions of the dataset, denoted as ScanNet3R-MSLAM and ScanNet3R-VGGT, respectively. Moreover, we leverage DINO-X to automatically produce view-wise open-vocabulary annotations for ScanNet3Rs as we have described in Sec 3.1.
@@ 4.2 Evaluation Settings
Open Setting. In this setting, the training data and annotations are not restricted. For example, some methods (Yang et al., 2023) are training free, some (Takmaz et al., 2023; Cao et al., 2025) are directly trained on the ScanNet200. For fair comparison, in this setting, we mix ScanNet200 and ScanNet3Rs for training. Following previous works, we report mAP25, mAP50, and mAP as the evaluation metrics. Specifically, mAP25 and mAP50 denote the mean Average Precision when the mask Intersection-over-Union (IoU) threshold is set to 25% and 50%, respectively, while mAP represents the average over multiple IoU thresholds ranging from 50% to 95% at a step of 5%. Moreover, to further analyze the comparison, the 200 classes are divided into head, common, and tail subsets according to their occurrence frequency, from high to low. We report the performance on the head and tail subsets separately, highlighting the class generalization ability among models.
@@ 4.4 Comparison with State-of-The-Arts
Open Setting. To quantify the gains from the additional training data and annotations provided by OVSeg3R, after extending SegDINO3D to SegDINO3D-VL, we first train it under the traditional training scheme, using only ScanNet200. Since the annotations at this setting cover only 200 categories, the resulting model remains closed-vocabulary. As shown in Table 1, under this setting, we observe a performance drop. We attribute this to the additional model capacity required for aligning text and visual features, which can also be observed in 2D models (Liu et al., 2024b). Moreover, although the model in principle supports open-vocabulary generalization, the limited diversity of training data constrains its semantic generalization ability. Consequently, its performance on tail classes remains significantly lower than that on head classes. As a comparison, as shown in the last row of Table 1, when SegDINO3D-VL is trained under the proposed OVSeg3R training scheme with open-vocabulary annotations, the model not only gains open-vocabulary capability but also achieves a notable overall improvement (+2.3 mAP), outperforming even the closed-vocabulary methods. The detailed performance on the tail and head categories reveals the source of the improvement. Benefiting from stronger class generalization ability, the model achieves a significant gain on tail classes (+8.7 mAP), reducing the performance gap with head classes from -11.3 mAP to -1.9 mAP.

### v2

@@ 1 Introduction
Recent advances in 3D reconstruction (mast3rslam; dust3r; vggt) have made scene geometry capturing accessible. Yet downstream tasks such as manipulation (liu2024rdt; black2024pi_0), navigation (song2025towards), and augmented reality (AR) require recognizing objects with instance-level identities and locations. Such demand has driven a growing interest in 3D instance segmentation and its open-vocabulary generalization.
Despite major progress in 2D segmentation (li2023mask), where open-vocabulary capabilities already meet most downstream demands (DINOX), 3D instance segmentation capabilities remain limited. This limitation persists even though the cost of acquiring 3D scenes has been greatly reduced by the remarkable progress of 3D reconstruction techniques, as the cost of acquiring 3D annotations is still expensive.
Therefore, how to leverage diverse 3D scenes provided by 3D reconstruction models and robust 2D masks provided by well-studied 2D perception models to enhance 3D instance segmentation has become an important research topic. Some approaches (openmask3d) train 3D segmentation models solely to produce class-agnostic 3D masks, which are then projected onto 2D to retrieve category information from 2D foundation models (clip). While this strategy can effectively exploit the strong classification ability of 2D models, it remains limited by the scarcity of 3D annotations, making it difficult to generate reliable 3D masks for unseen objects and often resulting in missed detections. Other methods (SAM3D) project each view’s 2D segmentation results (SAM) into 3D space using the 2D pixel to 3D point correspondences provided by a 3D reconstruction model, and merge the projected masks that belong to the same instance through heuristic strategies. Although this line of work benefits from the classification and segmentation strengths of 2D models, the hand-crafted merging process is error-prone, rendering these methods fragile and performance-constrained. While these methods demonstrate open-vocabulary potential, they over-rely on 2D outputs, leaving native 3D perception ability underdeveloped, which is crucial for advancing 3D understanding (openscene; ding2023pla). Although some works try to distill from 2D models, they need to train 3D Gaussian (kerbl20233d) to create 2D and 3D correspondences, which is redundant since reconstruction inherently provides this. Moreover, these methods either require per-scene optimization (lyu2024gaga; ye2024gaussian) or high-quality point cloud for Gaussian initialization (cao2025lift), limiting their practicality.
@@
To address the 3D annotation challenges, in this work, we propose a novel training scheme, called OVSeg3R. As shown in Fig 1, instead of relying on manually adjusted 3D scenes from non-routine sensors, OVSeg3R directly leverages a modern 3D reconstruction model such as (mast3rslam) or (vggt) to provide point cloud inputs. This not only substantially reduces the cost of acquiring 3D scenes, but also naturally introduces noise to the inputs, aligning with application scenarios, where user-provided inputs are typically videos or low-quality reconstructions captured by handheld sensors. For annotations, instead of manually annotating 3D masks for the reconstructed scene, the inherent 2D pixel-to-3D point correspondence provided by 3D reconstruction enables us to lift the 2D masks generated by an open-vocabulary 2D segmention model (DINOX) from each view into 3D space to obtain view-level annotations. However, although 2D masks from different views may correspond to the same object, they are estimated independently and lack cross-view associations (SAM3D). Directly concatenating them to a scene-level annotation would introduce many duplicate annotations. Conversely, since each view covers only part of the reconstructed point cloud, supervising scene-level predictions with each view’s partial annotations alone would incorrectly penalize predictions outside the view. To mitigate this issue, we propose a View-wise Instance Partition (VIP) algorithm. According to the visibility of each scene-level mask prediction’s corresponding object query in different views, VIP assigns the mask predictions to their corresponding views. Then, for each mask prediction, VIP further truncates it to retain only the region visible within its belonging view. The resulting view-level predictions enable supervision with view-specific annotations, which eliminates annotation duplication and avoids incorrect penalization, thus improving training stability.
@@
Moreover, to improve efficiency, mainstream 3D instance segmentation models (oneformer3d; segdino3d) normally leverage superpoints (landrieu2018large), first over-segmenting the input point cloud into superpoints and then performing instance segmentation at the superpoint level. However, in OVSeg3R, 3D reconstruction results are often over-smoothed (yang2024depth), leading to the loss of geometric details. As a result, constructing superpoints purely based on geometric continuity, as in the previous method, may cause objects that are not geometrically salient, such as paintings, being included in a superpoint covering large planar regions, such as walls, as shown in Fig 3 (b). This will inevitably lead to inaccurate segmentation. To mitigate this issue, we propose the 2D Instance Boundary-aware Superpoint (IBSp), IBSp incorporates 2D instance masks into the construction of superpoints, avoids the erroneous clustering of points from different instances into the same superpoint, further stabilizes the training process.
@@ 2 Related Works
Closed-vocabulary 3D Instance Segmentation. Early 3D instance segmentation methods fall into two lines: proposal-based methods (yang2019learning; hou20193d; yi2019gspn; engelmann20203d; td3d), which first detect objects and then refine 3D masks within the predicted bounding boxes, and grouping-based methods (liang2021instance; chen2021hierarchical; vu2022softgroup; jiang2020pointgroup; wang2019associatively; jiang2020end; zhang2021point), which aggregate points via voting in feature or geometric space. Following the success of Detection Transformers (DETR) (detr; liu2022dab; li2022dn; zhang2022dino) in 2D, recent works (spformer; Mask3D; maft; oneformer3d; ODIN) adopt DETR-like architectures for 3D instance detection and segmentation. Notably, SegDINO3D (segdino3d) proposes to leverage high-quality image- and object-level features from well-studied 2D models (DINOX; groundingdino) to support data-hungry 3D models, achieving substantial performance gains.
Open-vocabulary 3D Instance Segmentation. Motivated by the rapid progress in 2D open-vocabulary perception, most methods attempt to obtain open-vocabulary 3D perception outputs by relying on the outputs of 2D models. Generally, the mainstream methods can be categorized into two types. OpenMASK3D (openmask3d) first proposes to generate class-agnostic 3D segmentation results using 3D models (Mask3D). Then project each 3D result to 2D to obtain the corresponding object’s category using 2D foundation models (clip), thereby constructing an open-vocabulary 3D segmentor. While this approach achieves remarkable results and has inspired a series of subsequent works (open3dis; openyolo3d; nguyen2025any3dis), it only provides classification capability for novel categories. Limited by the scarcity of 3D segmentation data, such methods often fail to segment objects that are unseen during training. SAM3D (SAM3D) is the first to lift 2D segmentation results from SAM for each frame to 3D and then merge those belonging to the same instance in 3D using a proposed heuristic algorithm. Motivated by the promising performance, many subsequent works (SAI3D; ovir3d; zhao2025sam2object; xu2025sampro3d) have focused on improving this merging process. While these algorithms can reuse both the detection and classification capabilities of well-studied 2D models, the heuristics are fundamentally rule-based, lack generalization, and fail to handle numerous corner cases encountered in practical scenarios.
@@ 3 Method
To validate the effectiveness of the proposed training scheme, we adopt SegDINO3D (segdino3d), a recent state-of-the-art method, as our baseline. To further satisfy the requirement of open-vocabulary, we extend the classification part of SegDINO3D to the similarity calculation between object features and the text features, yielding SegDINO3D-VL. In this section, we will first describe the training of SegDINO3D-VL with OVSeg3R, including the preparation of data, construction of view-wise annotation, and obtaining predictions from SegDINO3D-VL. After that, we will describe in detail our designs for stable training.
@@ 3.1.1 Foundation Preparation
Given a video with $V$ views $\mathbf{I}\in\mathbb{R}^{V\times H\times W\times 3}$, where $H$ and $W$ are height and width, OVSeg3R feeds it to a 3D reconstructor and a 2D instance segmentator, instantiated as MASt3R-SLAM (mast3rslam) and DINO-X (DINOX) respectively by default, to prepare the foundation data.
@@ 3.1.2 Input and View-wise Annotation Preparation
2D Feature for Each 3D Point. Following SegDINO3D (segdino3d), we need to sample 2D image-level feature $\mathbf{F^{2D}}\in\mathbb{R}^{N\times C}$ for each 3D point. However, since the 2D-3D correspondences $\mathcal{C}$ are already available, re-projecting every 3D point to all views to identify a representative view, as required in SegDINO3D, becomes unnecessary. As described in ① of Fig 2, based on each 3D point’s corresponding view index and the sampling location provided by $\mathcal{C}$, we sample 2D features from 2D image feature maps $\mathbf{X}$ for the point through bilinear interpolation. For the sampling of $\mathbf{F}^{\mathbf{2D}}_{i}\in\mathbb{R}^{C}$ for the $i$-th 3D point, the process can be formulated as
@@
Text Prompt Feature. Following previous methods (groundingdino), we need to prepare the text features for each class name in $s$ for instance classification (③ in Fig 2). To enable contrastive learning and batch-friendly training, we randomly sample T|s|T-|s| additional class names not present in $s$ as negative classes, padding $s$ to a fixed size $T$. The padded $s$ is then concatenated into a string (e.g., ‘book . sofa .’) and encoded with a text encoder, to produce text features $\mathbf{T}\in\mathbb{R}^{T\times C}$.
@@ 3.2 View-wise Instance Partition
Compatibility with Scene-level 3D Instance Segmentation Supervision. It is worth noting that VIP is fully compatible with standard 3D instance segmentation supervision. For datasets that already provide scene-level annotations, there is no need to partition predictions to individual views for supervision; we can directly supervise the scene-level predictions using the scene-level annotations. This compatibility allows us to mix traditional costly annotated datasets, which provide manually adjusted reconstructed scenes and corresponding human annotations, with datasets that only have the raw videos, thereby enhancing the flexibility of the OVSeg3R training scheme. Moreover, for scene-level supervision, instead of performing a global matching between predictions and annotations via Hungarian matching (detr), we follow the previous method (oneformer3d; segdino3d) to sparsify the matching based on the relationship between the superpoints used to initialize object queries and the ground-truth masks. Specifically, a prediction can be matched to a ground-truth annotation only if the superpoint used to initialize its object query lies within that annotation’s mask. This design strengthens the connection between the superpoint used for object query initialization and the final mask prediction, thereby consolidating the foundation of VIP from the perspective of supervision and improving its reliability.
@@
Analysis. Current methods typically employ the Felzenszwalb (felzenszwalb2004efficient), a graph-based segmentation algorithm, to over-segment large point sets into compact superpoints, reducing the following computational overhead. The algorithm constructs a graph by connecting each 3D point to its K-Nearest Neighbors (KNN). Subsequently, adjacent points in the graph that satisfy geometric continuity constraints will be clustered into the same superpoint.
@@ 4.1 Datasets
We validate the effectiveness of OVSeg3R based on ScanNetv2 (scannet) and ScanNet200 (scannet200). They share the same 1,513 scenes, with 1,201 used for training and the remaining 312 for evaluation. The difference is that ScanNetv2 provides human-annotated instance masks for only 20 classes, whereas ScanNet200 extends the annotations to 200 classes. For each scene, both manually refined high-quality point clouds and the corresponding raw RGB videos are available. We reconstruct the raw videos using MASt3R-SLAM (mast3rslam) and VGGT (vggt), resulting in reconstructed versions of the dataset, denoted as ScanNet3R-MSLAM and ScanNet3R-VGGT, respectively. Moreover, we leverage DINO-X to automatically produce view-wise open-vocabulary annotations for ScanNet3Rs as we have described in Sec 3.1.
@@ 4.2 Evaluation Settings
Open Setting. In this setting, the training data and annotations are not restricted. For example, some methods (SAM3D) are training free, some (openmask3d; cao2025lift) are directly trained on the ScanNet200. For fair comparison, in this setting, we mix ScanNet200 and ScanNet3Rs for training. Following previous works, we report mAP25, mAP50, and mAP as the evaluation metrics. Specifically, mAP25 and mAP50 denote the mean Average Precision when the mask Intersection-over-Union (IoU) threshold is set to 25% and 50%, respectively, while mAP represents the average over multiple IoU thresholds ranging from 50% to 95% at a step of 5%. Moreover, to further analyze the comparison, the 200 classes are divided into head, common, and tail subsets according to their occurrence frequency, from high to low. We report the performance on the head and tail subsets separately, highlighting the class generalization ability among models.
@@ 4.4 Comparison with State-of-The-Arts
Open Setting. To quantify the gains from the additional training data and annotations provided by OVSeg3R, after extending SegDINO3D to SegDINO3D-VL, we first train it under the traditional training scheme, using only ScanNet200. Since the annotations at this setting cover only 200 categories, the resulting model remains closed-vocabulary. As shown in Table 1, under this setting, we observe a performance drop. We attribute this to the additional model capacity required for aligning text and visual features, which can also be observed in 2D models (groundingdino). Moreover, although the model in principle supports open-vocabulary generalization, the limited diversity of training data constrains its semantic generalization ability. Consequently, its performance on tail classes remains significantly lower than that on head classes. As a comparison, as shown in the last row of Table 1, when SegDINO3D-VL is trained under the proposed OVSeg3R training scheme with open-vocabulary annotations, the model not only gains open-vocabulary capability but also achieves a notable overall improvement (+2.3 mAP), outperforming even the closed-vocabulary methods. The detailed performance on the tail and head categories reveals the source of the improvement. Benefiting from stronger class generalization ability, the model achieves a significant gain on tail classes (+8.7 mAP), reducing the performance gap with head classes from -11.3 mAP to -1.9 mAP.